OpenStudy (anonymous):

what is 3c + 4 = -20?

Parth (parthkohli):

``` what is 3c + 4 = -20 ``` it is an equation.

OpenStudy (anonymous):

lol mis read Q

OpenStudy (triciaal):

if you need to solve for c subtract 4 from each side to leave the variable on one side and the numbers on the other 3c + 4 -4 = -20-4 3c = -24 divide each side by 3 because you have 3*c and you want 1*c whatever you do to one side you do to the other to keep the balance of the eqn 3c/3 = =24/3 c = -8